OpenStudy (anonymous):

what is the simplified form of 1/x-2/x^2+x ?

OpenStudy (luigi0210):

Welcome to Openstudy @emilyporazinski :)

OpenStudy (luigi0210):

Is this your question? \[\frac{\frac{1}{x-2}}{x^2+x}\]

OpenStudy (anonymous):

No its 1/x - 2/x^2+x.....if that helps see it better

OpenStudy (luigi0210):

\[\dfrac{1}{x}-\frac{2}{x^2+x}\]

OpenStudy (anonymous):

yes that itttttt

OpenStudy (anonymous):

take LCM and then divide it.

OpenStudy (luigi0210):

Okay, so what you have to do is get common denominators

OpenStudy (anonymous):

ohhhhhhh! thank you(:

OpenStudy (luigi0210):

Yup :)

OpenStudy (anonymous):

you get it @emilyporazinski

OpenStudy (anonymous):

waiiiiiit, what would the least common denominator be??

OpenStudy (luigi0210):

Well one has x, and another has x^2+x, so it should be x^2+x

OpenStudy (anonymous):

yes @Luigi0210 is right LCM is x2+x